OpenStudy (anonymous):

Express sin 186 degrees as a sine function of a positive acute angle

OpenStudy (anonymous):

not sure how you are going to get an acute angle out of this it will be the same as \(\sin(-9)\) perhaps you can say it is \(-\sin(9)\)

OpenStudy (anonymous):

how though?

OpenStudy (anonymous):

scratch that, i am not paying attention it is 186 degrees not 189

OpenStudy (anonymous):

so it is the same as \(-\sin(6)\)

OpenStudy (anonymous):

let me draw a picture

OpenStudy (anonymous):



OpenStudy (anonymous):

since \(186\) is in quadrant 3, you are below the horizontal axis and so sine is negative go across the unit circle to get to 6 degrees, and acute angle, that has the same value as \(\sin(186\) except it is positive so \[\sin(186)=-\sin(6)\] don't let the minus sign confuse you \(\sin(186)\) is negative so \(-\sin(9)\) is positive

OpenStudy (anonymous):

allrightt i got it thank you!!